The enumerate package^{∗}

David Carlisle carlisle@cs.man.ac.uk

1994/01/31

Abstract

This package gives the enumerate environment an optional argument which determines the style in which the counter is printed.

An occurence of one of the tokens A a I i or 1 produces the value of the counter printed with (respectively) \Alph \alph \Roman \roman or

\arabic.

These letters may be surrounded by any strings involving any other TEX expressions, however the tokens A a I i 1 must be inside a { } group if they are not to be taken as special.

1 Examples

EX i. one one one one one one one one one one one one

EX ii. two

example a) one of two one of two one of two

example b) two of two

A-1 one

A-2 two

\begin{enumerate}[EX i.]
\item one one one one one one one
       one one one one\label{LA}
\item two
    \begin{enumerate}[{example} a)]
    \item one of two  one of two
          one of two\label{LB}
    \item two of two
    \end{enumerate}
    \item two of two
    \end{enumerate}
\end{enumerate}
\begin{enumerate}[{A}-1]
\item one\label{LC}
\item two
\end{enumerate}

\label and \ref may be used as with the standard enumerate environment. \ref only produces the counter value, not the whole label. \ref prints the value in the same style as \item, as determined by the presence of one of the tokens A a I i 1 in the optional argument. In the above example \ref{LA}, \ref{LB} and \ref{LC} produce ‘i’, ‘iia’ and ‘1’ respectively.

∗This file has version number v2.02, last revised 1994/01/31.